OpenStudy (anonymous):

can somebody help me with this? Identify the slope and y-intercept in the following equations. 2x+3y=8

OpenStudy (anonymous):

You must first rearrange the function into slope-intercept form;\[y=mx+b\]Can you do that?

OpenStudy (anonymous):

37=2x+8?

OpenStudy (anonymous):

I mean 3y

Mehek (mehek14):

You have to get rid of 3 in 3y

OpenStudy (anonymous):

solve for y now.

OpenStudy (dmndlife24):

no -2x

OpenStudy (dmndlife24):

signs change

OpenStudy (dmndlife24):

3y=-2x+8

OpenStudy (anonymous):

Almost. Except you subtracted 2x from both sides so there is a mistake with one of the signs. Can you see it?

OpenStudy (dmndlife24):

\[y = mx + b\] \[3y = -2x + 8\]

OpenStudy (anonymous):

so 3y= -2x + 8 is right?

OpenStudy (anonymous):

Correct. Now how do you solve for y?

OpenStudy (anonymous):

am I supposed to get the y by itself?

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

do I need to divide by 3/1

OpenStudy (anonymous):

That's right. Divide both sides by 3. What do you get?

OpenStudy (anonymous):

I cant figure it out I know it makes 3y = y but I don't know how to divide the other side by 3?

OpenStudy (anonymous):

OK. You're correct about the left side. Now, on the right side you have to remember to divide all the terms by 3. The first term is -2x. What do you get when you divide -2 by 3?

OpenStudy (anonymous):

-2/3x

OpenStudy (anonymous):

Great. I assume you mean\[-\frac{ 2 }{ 3 }x\]The seond term on the right side is 8. What do you get when you divide 8 by 3?

OpenStudy (anonymous):

2 2/3

OpenStudy (anonymous):

Terrific. Now, to sum up you have\[y=-\frac{ 2 }{ 3 }x + 2\frac{ 2 }{ 3 }\]That's the slope-intercept form. Compare this with\[y=mx+b\]Can you determine the value of m and b?

OpenStudy (anonymous):

m is the coefficient of x, and b is the constant on the end.

OpenStudy (anonymous):

yes, so the slope is 2/3 and the y intercept is 2 2/3?

OpenStudy (anonymous):

Oh, so close. You've made an error with a sign. See it?

OpenStudy (anonymous):

slope = -2/3?

OpenStudy (anonymous):

Terrific. Great job! Your other questions should proceed in much the same way.

OpenStudy (anonymous):

thank you very much!